OpenStudy (anonymous):

Write a program that prompts the user for a Celsius temperature. Your program will calculate and print the Fahrenheit equivalent

OpenStudy (bibby):

c->f Multiply by 9, then divide by 5, then add 32 \(\huge \frac{9C}{5}+32\)

OpenStudy (mandre):

1. Celsius = user input. 2. Fahrenheit = (9/5)*Celsius + 32 3. Output(Fahrenheit). Now all you need to do is convert this to your programming language of choice.

OpenStudy (arnavguddu):

C/5 = (F-32)/9 Therefore, F = 9*C/5+32